# A primer on quasi-convex functions in nonlinear potential theories

### v2

@@ Acknowledgements
    and
    $\mathscr{L}g=\mathscr{L}\!\mathscr{L}f=|\cdot|\quad\text{on}\quad\partial g(\partial f(X))=\partial g([-1,1])={\mathbb{R}}\supsetneqq(-1,1)=X.$
    For the sake of completeness, we also mention that this example is based on the fact that the function $\tilde{f}$ defined on ${\mathbb{R}}$ by $\tilde{f}=f$ in $(-1,1)$ and $f\equiv+\infty$ in ${\mathbb{R}}\setminus(-1,1)$ is not a closed convex function. On the other hand, the Fenchel–Moreau Theorem tells that if $f\colon{\mathbb{R}}^{n}\to{\mathbb{R}}$ is defined on the whole ${\mathbb{R}}^{n}$ (that is, we have $f\colon{\mathbb{R}}^{n}\to{\mathbb{R}}$ convex), then the equality $\mathscr{L}\!\mathscr{L}f=f$ is equivalent to $f$ being closed. Since one can always think of $f\colon X\to{\mathbb{R}}$ as $\tilde{f}\colon{\mathbb{R}}^{n}\to{\mathbb{R}}$ letting $\tilde{f}\equiv+\infty$ on ${\mathbb{R}}^{n}\setminus X$, in such a way that $\partial f(x)=\partial\tilde{f}(x)$ for each $x\in X$ while $\partial\tilde{f}(x)=\emptyset$ if $x\notin X$ in order not to alter the domain $\partial f(X)$ of the Legendre transform, one sees that $\mathscr{L}\!\mathscr{L}f=f$ if and only if $\mathrm{dom}\,\mathscr{L}\!\mathscr{L}\tilde{f}\vcentcolon=\{x\in{\mathbb{R}}^{n}:\ \mathscr{L}\!\mathscr{L}\tilde{f}<+\infty\}=\mathrm{dom}\,\tilde{f}=X$, it and only if $\mathscr{L}\!\mathscr{L}\tilde{f}=\tilde{f}$, if and only if $\tilde{f}$ is closed, which forces $X={\mathbb{R}}^{n}$. Therefore, we conclude that one has $\mathscr{L}\!\mathscr{L}f=f$ in Proposition A.3(iv) only in the trivial case $X={\mathbb{R}}^{n}$. The reader can consult [42] or [48] for further details.
    Having defined and examined some basic properties of the Legendre transform of an arbitrary convex function, we will know show that much more can said about the Legendre transform of convex functions of the form
@@
      the subdifferential $\partial f\colon X\to\mathscr{P}({\mathbb{R}}^{n})$ is an expansive map; i.e.,
@@
      the subdifferential $G\vcentcolon=\partial g$ of $g=\mathscr{L}f$ is a single-valued function from $\partial f(X)$ onto $X$, and it is the inverse of $\partial f$; that is,
@@
    If, in addition, $u$ is bounded with $M\vcentcolon=\sup_{X}|u|$, $\delta\vcentcolon=2\sqrt{rM}$ and $X_{\delta}\vcentcolon=\{x\in X:d(x,\partial X)>\delta\}$, then the following hold:
@@
      $X_{\delta}\subset\partial f(X)$ and $\partial f(X_{\delta})\subset X$;
@@
      $g\vcentcolon=\mathscr{L}f\in C^{1}(X_{\delta})$, with $Dg=G$ on $X_{\delta}$.
@@
    $\langle p_{2}-p_{1},x_{2}-x_{1}\rangle\geqslant 0,\quad\forall\,p_{j}\in\partial u(x_{j}),j=1,2,$ (A.11)
@@
    For the claims of part (b), let $x_{1},x_{2}\in G(y)\cap X$ with $y\in\partial f(X)$. Then by Proposition A.3(c) one has $y\in\partial f(x_{1})\cap\partial f(x_{2})$, and thus $x_{1}=x_{2}$ by (A.8). Therefore $G(\partial f(x))\cap X=\{x\}$ for each $x\in X$, and in fact $G(\partial f(x))=\{x\}$ for each $x\in X$; indeed, this easily follows from the fact that $G(\partial f(x))$ is convex (by Lemma 1.10(a) and Proposition A.3(b)) and $X$ is open. This shows that $G$ is single-valued and it is the inverse of $\partial f$.
@@
    Now assuming that $u$ convex is also bounded, we verify the claims of parts (d) and (e). First for $y\in X_{\delta}$ we will show that $y\in\partial f(X)$. To that end, we first note that the continuous function defined by $h_{y}\vcentcolon=ru+\frac{1}{2}|\cdot-\,y|^{2}$ satisfies
@@
    $h_{y}(z)-h_{y}(y)=ru(z)+\tfrac{1}{2}|z-y|^{2}-ru(y)>-2rM+\tfrac{1}{2}\delta^{2}=0,$
    therefore $\min_{\overline{B}_{\delta}(y)}h_{y}\leqslant h_{y}(y)\leqslant\inf_{X\setminus\overline{B}_{\delta}(y)}h_{y}$. Now let $x\in\overline{B}_{\delta}(y)$ be a point which realizes the minimum of $h_{y}$ on $X$. Hence, $0\in\partial h_{y}(x)$; that is, $0=rp+(x-y)$ for some $p\in\partial u(x)$. Hence, we have that $y\in\partial f(X)$ by (A.10), which proves the first inclusion in part (d).
    Next, for $y\in\partial f(X_{\delta})$ we will show that $x\in X$. Since $y\in\partial f(X_{\delta})$, we have $y\in\partial f(x)$ for some $x\in X_{\delta}$. We know that $f-\langle y,\cdot\rangle$ has a minimum point on $X$ at $x$. This is equivalent to $h_{y}$ having a minimum point on $X$ at $x$, since by definition $h_{y}=f-\langle y,\cdot\rangle+\frac{1}{2}|y|^{2}$. Then, $h_{y}$ has a minimum point at $x$, and we have
@@
    where we used (A.12) for the last inequality, since $|y-w|\geqslant d(y,\overline{X})=d(y,X)$. This contradicts the hypothesis that $x\in X_{\delta}$, thus $y\in X$.
@@
    The proof of Alexandrov’s theorem will depend on two ingredients (Lemma A.7 and Theorem A.9 below). The former is a crucial property of the Legendre transform, which provides a useful sufficient condition for $f$ defined by (A.7), with $u$ convex and bounded, to be twice differentiable. It essentially states that the subdifferential $G$ of the Legendre transform $g$ of $f$ maps the complement of its critical points $k_{G}$ to a set of points where $f$ is twice differentiable; that is, $G(X_{\delta}\setminus k_{G})\subset\mathrm{Diff}^{2}f$, where $X_{\delta}$ is defined as in Proposition A.5. The latter is a Lipschitz version of Sard’s theorem. It allows one to handle the set of “bad” points (namely the set of critical values of $G$, which is a Lipschitz function by Proposition A.5(c)), by showing that this set has zero Lebesgue measure.
    Let us recall what critical points and critical values are, state and prove the former ingredient (Lemma A.7). Then we will state the latter ingredient (Theorem A.9) and use it in the proof of Alexandrov’s theorem, while the reader is invited to have a look at the next Appendix LABEL:proofsard for a technical measure-theoretic proof of Theorem A.9, based on Besicovitch’s covering theorem.
@@
    The fist needed ingredient is the following.
    Lemma A.7 (Magic property of the Legendre transform).
    Let $f$ be defined by (A.7); that is, $f:=ru+\frac{1}{2}|\cdot|^{2}$ with $r>0$ a real number and $u$ a convex function which is bounded on $X$. Let $G=\partial g=\partial\mathscr{L}f$ and $X_{\delta}$ be defined as in Proposition A.5. Suppose that
@@
    In short, Lemma A.7 tells that, if $|u|\leqslant M$ on $X$, $f\vcentcolon=ru+\frac{1}{2}|\cdot|^{2}$ with $r>0$, and $G=\partial\mathscr{L}f$, then
    $G\!\left(X_{2\sqrt{rM}}\setminus k_{G}\right)\cap\mathrm{Diff}^{1}f\subset\mathrm{Diff}^{2}f$
@@
    $D^{2}u\circ G=r^{-1}((DG)^{-1}-I)\quad\text{on $\big{(}X_{2\sqrt{rM}}\setminus k_{G}\big{)}\cap G^{-1}(\mathrm{Diff}^{1}f)$}.$
@@
    By the hypotheses (i) and (ii), 00 is not a critical value of $G$ and hence $B:=DG(0)$ is invertible. The proof of the lemma reduces to showing that
@@
    By Proposition A.5(d), given $x\in X_{\delta}$ there exists $y\in X$ such that $y\in\partial f(x)$, then by Proposition A.3(c), $x\in\partial g(y)=\{G(y)\}$; that is, $x=G(y)$. For this pair $(x,y)\in\partial f$, by the definition of the Legendre transform one can write
    $f(x)-Q_{A}(x)=Q_{B}(y)-g(y)+\tfrac{1}{2}\langle y-Ax,x\rangle+\tfrac{1}{2}\langle x-By,y\rangle.$ (A.14)
    Since $G$ is differentiable in $y_{0}=0$ with $G(0)=0$ and $B=DG(0)$, one has
@@
    The asymptotic expansion (A.15) allows us to show that the Legendre transform $g$ of $f$ is twice differentiable with
@@
    By definition $g(0)=0$ (since $f(0)=0$ and $(0,0)\in\partial f$) and by hypothesis $G(0)=Dg(0)=0$. Hence the claim (A.17) is equivalent to
@@
    Indeed, applying the mean value theorem to the function $h:=g-Q_{B}$ which satisfies $h(0)=0$ there exists $z\in[0,y]$ such that $h(y)=\langle Dh(z),y\rangle$, which is (A.19). Using (A.15) one has
@@
    Finally, we conclude the proof of the lemma by establishing (A.13). Using (A.15), (A.16) and (A.18) on the terms on the right-hand side of (A.14) yields
    $f(x)-Q_{A}(x)=o(|y|^{2})\quad\text{for $y\to 0$},$ (A.20)
@@
    $|y|=|ABy|\leqslant||A||\,(|x-By|+|x|),\quad\forall\,x,y\in{\mathbb{R}}^{n},$ (A.22)
@@
    The second needed ingredient is the following, which will be proven in Appendix LABEL:proofsard.
@@
    Let $\Omega\subset{\mathbb{R}}^{n}$ be open. The set of critical values of a Lipschitz function $G:\Omega\subset{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$ has (Lebesgue) measure zero.
@@
    Without loss of generality, we can suppose that $u$ is in fact bounded on $X$; indeed, if not, it suffices to consider an exhaustion by compact convex sets $K_{j}\nearrow X$ (for instance, one can take $K_{j}\vcentcolon=\big{\{}x\in X:\ d(x,\partial X)\geqslant 2^{-j},\,|x|<2^{j}\big{\}}$ for all $j\in\mathbb{N}$), prove the theorem for $u|_{\operatorname{int}\!K_{j}}$ (which is bounded since $u$ is continuous on $X$; cf. Theorem 1.14), and then note that the set of all points in $X$ where $u$ is twice differentiable, $\mathrm{Diff}^{2}u$, has full measure as well, since
@@
    which, by the $\sigma$-subadditivity of the Lebesgue measure, yields $|X\setminus\mathrm{Diff}^{2}u|=0$.
    Hence, we will prove the theorem for $u$ bounded on $X$. Fix $r>0$ and let $f\vcentcolon=ru+\frac{1}{2}|\cdot|^{2}$. By Lemma A.7 (cf. also Remark A.8), we know that
    $G(X_{\delta}\setminus k_{G})\cap\mathrm{Diff}^{1}u\subset\mathrm{Diff}^{2}u,$ (A.23)
    where (cf. Proposition A.5)
@@
    $\big{|}\mathrm{Diff}^{2}u\big{|}\geqslant\big{|}G(X_{\delta})\big{|}\geqslant\big{|}X_{2\delta}\big{|},$ (A.25)
    where the latter inequality comes from the fact that $X_{2\delta}\subset G(X_{\delta})$. Indeed, by Proposition A.5(d), $\partial f(X_{2\delta})\subset X_{\delta}$, which by part (b) of the same Proposition is equivalent to $X_{2\delta}\subset G(X_{\delta})$, as we claimed.
    Notice now that $r>0$ is arbitrary, thus we can let $r\searrow 0$; that is, (cf. (A.24)) we can let $\delta\searrow 0$ in (A.25). We obtain
    $\big{|}\mathrm{Diff}^{2}u\big{|}\geqslant|X\big{|},$
    then yields $\big{|}\mathrm{Diff}^{2}u\big{|}=|X\big{|}$, since the reverse inequality holds trivially. This proves that $\mathrm{Diff}^{2}u$ has full measure (in $X$), which is the desired conclusion. ∎
    Appendix B.
    \@afterheading
    app:bes Let $A\subset{\mathbb{R}}^{n}$ and $\rho\,\colon A\to{\mathbb{R}}_{+}$ bounded. Then there exists $N\in\mathbb{N}$ depending only on $n$ and $N$ families

### v3

@@ Appendix A The Legendre transform and Alexandrov’s theorem
and $\mathscr{L}g=|\cdot|$ on $\partial g(\partial f(X))=\partial g([-1,1])={\mathbb{R}}\supsetneqq X$.
For the sake of completeness, we mention that this example is based on the fact that the function $\tilde{f}$ defined on ${\mathbb{R}}$ by $\tilde{f}=f$ in $(-1,1)$ and $f\equiv+\infty$ in ${\mathbb{R}}\setminus(-1,1)$ is not lower semicontinuous function. Indeed, the Fenchel–Moreau Theorem states that given $f\colon{\mathbb{R}}^{n}\to(-\infty,+\infty]$ proper (that is, $\mathrm{dom}(f)\vcentcolon=\{x\in{\mathbb{R}}^{n}:\,f(x)<+\infty\}\neq\emptyset$), the identity $\mathscr{L}\!\mathscr{L}f=f$ is equivalent to $f$ being lower semicontinuous and convex. In other words, part (d) of Proposition A.4 simply becomes the identity $\mathscr{L}\mathscr{L}f=f$ (which also entails that $\partial g(\partial f(X))=X$) when $f\colon X\to{\mathbb{R}}$ is convex and continuous on $X$ convex and closed.
On the other hand, in general, the double transform $\mathscr{L}\mathscr{L}f$ provides the largest lower semicontinuous convex function $g\leqslant f$ (as it can be noted in the above example). The reader can consult, e.g., [47, 51] for further details.
Having defined and examined some basic properties of the Legendre transform of an arbitrary function, we will know show that much more can said about the Legendre transform of convex functions of the form
@@
    the subdifferential $\partial f\colon X\to\mathscr{P}({\mathbb{R}}^{n})$ is an expansive map; that is,
@@
    the subdifferential $G\vcentcolon=\partial g$ of $g\vcentcolon=\mathscr{L}f$ is a single-valued function from $\partial f(X)$ onto $X$, and it is the inverse of $\partial f$; that is,
@@
If, in addition, $u$ is bounded, then, letting $M\vcentcolon=\sup_{X}|u|$, $\delta\vcentcolon=2\sqrt{rM}$ and $X^{\delta}\vcentcolon=\{x\in X:d(x,\partial X)>\delta\}$, the following hold:
@@
    $X^{\delta}\subset\partial f(X)$ and $\partial f(X^{\delta})\subset X$;
@@
    $g\vcentcolon=\mathscr{L}f\in C^{1}(X^{\delta})$, with $Dg=G$ on $X^{\delta}$.
@@
$\langle p_{2}-p_{1},x_{2}-x_{1}\rangle\geqslant 0,\quad\forall\,p_{j}\in\partial u(x_{j}),\ j=1,2,$
@@
For the claims of part (b), let $x_{1},x_{2}\in G(y)\cap X$ with $y\in\partial f(X)$. Then by Proposition A.4(c) one has $y\in\partial f(x_{1})\cap\partial f(x_{2})$, and thus $x_{1}=x_{2}$ by (A.8). Therefore $G(\partial f(x))=\{x\}$ for each $x\in X$; indeed, this easily follows from the fact that $G(\partial f(x))$ is convex (by Lemma 1.12(a) and Proposition A.4(b)) and $X$ is open. This shows that $G$ is single-valued and it is the inverse of $\partial f$.
@@
Now assuming that $u$ convex is also bounded, we verify the claims of parts (d) and (e). First for $y\in X^{\delta}$ we will show that $y\in\partial f(X)$. To that end, we first note that the continuous function defined by $h_{y}\vcentcolon=ru+\frac{1}{2}|\cdot-\,y|^{2}$ satisfies
@@
$h_{y}(z)-h_{y}(y)=ru(z)+\tfrac{1}{2}|z-y|^{2}-ru(y)>-2rM+\tfrac{1}{2}\delta^{2}=0;$
therefore, $\min_{\overline{B}_{\delta}(y)}h_{y}\leqslant h_{y}(y)\leqslant\inf_{X\setminus\overline{B}_{\delta}(y)}h_{y}$. Now let $x\in\overline{B}_{\delta}(y)$ be a point which realizes the minimum of $h_{y}$ on $X$. Hence, $0\in\partial h_{y}(x)$; that is, $0=rp+(x-y)$ for some $p\in\partial u(x)$. We deduce that $y\in\partial f(X)$ by (A.10), which proves the first inclusion in part (d).
Next, for $y\in\partial f(X^{\delta})$ we will show that $x\in X$. Since $y\in\partial f(X^{\delta})$, we have $y\in\partial f(x)$ for some $x\in X^{\delta}$. We know that $f-\langle y,\cdot\rangle$ has a minimum point on $X$ at $x$. This is equivalent to $h_{y}$ having a minimum point on $X$ at $x$, since by definition $h_{y}=f-\langle y,\cdot\rangle+\frac{1}{2}|y|^{2}$. Then, $h_{y}$ has a minimum point at $x$, and we have
@@
where we used (A.11) for the last inequality, since $|y-w|\geqslant d(y,\overline{X})=d(y,X)$. This contradicts the hypothesis that $x\in X^{\delta}$, thus $y\in X$.
@@
As anticipated, the proof of Alexandrov’s theorem will depend on two ingredients (Lemma A.8 and Theorem A.10 below). The former is a crucial property of the Legendre transform, which provides a useful sufficient condition for $f$ defined by (A.7), with $u$ bounded, to be twice differentiable. It essentially states that the subdifferential $G$ of the Legendre transform $g$ of $f$ maps the complement of its critical points $k_{G}$ to a set of points where $f$ is twice differentiable; that is, $G(X^{\delta}\setminus k_{G})\subset\mathrm{Diff}^{2}f$, where $X^{\delta}$ is defined as in Proposition A.6. The latter is a Lipschitz version of Sard’s theorem. It allows one to handle the set of “bad” points (namely the set of critical values of $G$, which is a Lipschitz function by Proposition A.6(c)), by showing that this set has zero Lebesgue measure.
Let us recall what critical points and critical values are, and state and prove the former ingredient (Lemma A.8). Then we will state the latter ingredient (Theorem A.10) and use it in the proof of Alexandrov’s theorem, while the reader is invited to have a look at the next Appendix B for a technical measure-theoretic proof of Theorem A.10, based on Besicovitch’s covering theorem.
@@
The first needed ingredient is the following.
Lemma A.8 (Magical property of the Legendre transform).
Let $f$ be defined by (A.7); that is, $f\vcentcolon=ru+\frac{1}{2}|\cdot|^{2}$ with $r>0$ a real number and $u$ a convex function which is bounded on $X$. Let $G$ and $X^{\delta}$ be defined as in Proposition A.6. Assume the following:
@@
In short, Lemma A.8 tells that, if $|u|\leqslant M$ on $X$, $f\vcentcolon=ru+\frac{1}{2}|\cdot|^{2}$ with $r>0$, $G\vcentcolon=\partial\mathscr{L}f$, then, with $\delta\vcentcolon=2\sqrt{rM}$,
$G\bigl{(}X^{\delta}\setminus k_{G}\bigr{)}\cap\mathrm{Diff}^{1}f\subset\mathrm{Diff}^{2}f$
@@
$D^{2}u\circ G=r^{-1}((DG)^{-1}-I)\quad\text{on $\big{(}X^{\delta}\setminus k_{G}\big{)}\cap G^{-1}(\mathrm{Diff}^{1}f)$}.$
@@
By the hypotheses (i) and (ii), 00 is not a critical value of $G$ and hence $B\vcentcolon=DG(0)$ is invertible. The proof of the lemma reduces to showing that
@@
Step 1. By Proposition A.6(d), given $x\in X^{\delta}$ there exists $y\in X$ such that $y\in\partial f(x)$, then by Proposition A.4(c), $x\in\partial g(y)=\{G(y)\}$; that is, $x=G(y)$. For this pair $(x,y)\in\partial f$, by the definition of the Legendre transform one can write
.$f(x)-Q_{A}(x)=Q_{B}(y)-g(y)+\tfrac{1}{2}\langle y-Ax,x\rangle+\tfrac{1}{2}\langle x-By,y\rangle.$ (A.13)
Step 2. Since $G$ is differentiable in $y_{0}=0$ with $G(0)=0$ and $B=DG(0)$, one has
@@
Step 3. The asymptotic expansion (A.14) allows us to show that the Legendre transform $g$ of $f$ is twice differentiable with
@@
Note that by definition $g(0)=0$ (since $f(0)=0$ and $(0,0)\in\partial f$) and by hypothesis $G(0)=Dg(0)=0$. Hence the claim (A.16) is equivalent to
@@
Then, using (A.14) one has
@@
Step 4. Finally, we conclude the proof of the lemma by establishing (A.12). Using (A.14), (A.15) and (A.17) on the terms on the right-hand side of (A.13) yields
$f(x)-Q_{A}(x)=o(|y|^{2})\quad\text{for $y\to 0$},$
@@
$|y|=|ABy|\leqslant||A||\,(|x-By|+|x|),\quad\forall\,x,y\in{\mathbb{R}}^{n},$
@@
The second needed ingredient is the following, which will be proven in Appendix B.
@@
Let $\Omega\subset{\mathbb{R}}^{n}$ be open. The set of critical values of a Lipschitz function $G:\Omega\to{\mathbb{R}}^{n}$ has Lebesgue measure zero.
@@
Without loss of generality, we can suppose that $u$ is in fact bounded on $X$; indeed, if not, it suffices to consider an exhaustion by compact convex sets $K_{j}\nearrow X$ (for instance, one can take $K_{j}\vcentcolon=\big{\{}x\in X:\ d(x,\partial X)\geqslant 2^{-j},\,|x|<2^{j}\big{\}}$ for all $j\in\mathbb{N}$), prove the theorem for $u|_{\operatorname{int}\!K_{j}}$ (which is bounded since $u$ is continuous on $X$ by Theorem 1.16), and then note that $\mathrm{Diff}^{2}u$ has full measure as well, since
@@
which by the $\sigma$-subadditivity of the Lebesgue measure yields $|X\setminus\mathrm{Diff}^{2}u|=0$.
Hence, we will prove the theorem for $u$ bounded on $X$. Fix $r>0$ and let $f\vcentcolon=ru+\frac{1}{2}|\cdot|^{2}$. By Lemma A.8 (see also Remark A.9), we know that
$G(X^{\delta}\setminus k_{G})\cap\mathrm{Diff}^{1}u\subset\mathrm{Diff}^{2}u,$ (A.20)
where (see Proposition A.6)
@@
$\big{|}\mathrm{Diff}^{2}u\big{|}\geqslant\big{|}G(X_{\delta})\big{|}\geqslant\big{|}X^{2\delta}\big{|},$ (A.22)
where the latter inequality comes from the fact that $X^{2\delta}\subset G(X^{\delta})$. Indeed, by Proposition A.6(d), $\partial f(X^{2\delta})\subset X^{\delta}$, which by part (b) of the same proposition is equivalent to $X^{2\delta}\subset G(X^{\delta})$, as we claimed.
Notice now that $r>0$ is arbitrary, thus we can let $r\searrow 0$; that is, recalling (A.21), we can let $\delta\searrow 0$ in (A.22). We obtain $|\mathrm{Diff}^{2}u|\geqslant|X|$, and thus $|\mathrm{Diff}^{2}u|=|X|$, since the reverse inequality holds trivially. This proves that $\mathrm{Diff}^{2}u$ has full measure (in $X$), which is the desired conclusion. ∎
Appendix B Proof of Sard’s theorem for Lipschitz functions
Since Theorem A.10 states that a set of “bad” points (the critical values of a Lipschitz function $G$) has zero measure, it is not surprising that one needs to appeal to some measure-theoretic covering lemma. We will make use of a theorem of Besicovitch [5], which we will prove in Lemma B.2 by adapteing the proof of a more general form contained in [20].
We will also need the following elementary estimate on controlling the Lebesgue outer measure $m$ (see, e.g., [49, Chapter 1] for its definition and basic properties) of images of small balls about points of differentiability of a vector field $G$ on ${\mathbb{R}}^{n}$.
Lemma B.1.
Suppose that $G\colon\Omega\to{{\mathbb{R}}^{n}}$ is differentiable at $a\in\Omega$ with $\Omega\subset{\mathbb{R}}^{n}$ open. Then for each $\varepsilon>0$ there exists $\overline{\rho}=\overline{\rho}(a,\varepsilon)>0$ such that
$m\big{(}G\big{(}B_{\rho}(a)\big{)}\big{)}\leqslant\left(|{\det DG(a)}|+\varepsilon\right)m(B_{\rho}(a))\qquad\forall\rho\leqslant\overline{\rho}.$
Proof.
Since $G$ is differentiable at $a$, one has that
$|G(x)-\ell(x)|\leqslant\omega(|x-a|)|x-a|,\quad\text{for all $x$ near $a$},$
where :$\ell\vcentcolon=G(a)+\langle DG(a),\cdot-a\rangle$ is the linearization of $G$ at $a$ and $\omega\colon{\mathbb{R}}_{+}\to{\mathbb{R}}_{+}$ is a modulus of continuity (so $\omega(0^{+})=0$). This yields
$G\big{(}B_{\rho}(a)\big{)}\subset\ell\big{(}B_{\rho}(a)\big{)}+\omega(\rho)\rho B_{1}(0),$
By the monotonicity of the outer measure $m$, one has
$m(G(B_{\rho}(a)))\leqslant m(\ell(B_{1}(a))+\omega(\rho)B_{1}(0))\rho^{n},$
Now note that
$\lim_{\rho\searrow 0}m(\ell(B_{1}(a))+\omega(\rho)B_{1}(0))=m(\ell(B_{1}(a)))=|\ell(B_{1}(a))|=|{\det DG(a)}|\,\alpha_{n},$
where $\alpha_{n}\vcentcolon=m(B_{1}(0))$ and the latter equality comes from the change of variables formula. Therefore, there exists some $\overline{\rho}>0$ such that for $\rho\leqslant\overline{\rho}$ we have
$m\big{(}G\big{(}B_{\rho}(a)\big{)}\big{)}\leqslant m(\ell(B(a))+\omega(\rho)B)\rho^{n}\leqslant(|{\det DG(a)}|+\varepsilon)\,\alpha_{n}\rho^{n}$
and hence the result. ∎
Next is the needed covering lemma.
Lemma B.2 (Besicovitch’s covering theorem).
Let $A\subset{\mathbb{R}}^{n}$ and $\rho\,\colon A\to{\mathbb{R}}_{+}$ be bounded. Then there exists $N\in\mathbb{N}$, depending only on $n$, and $N$ families of balls